# Octopus v2: On-device language model for super agent

### v4

@@ 1 Introduction
Large language models have demonstrated impressive capabilities in function calling, significantly contributing to AI agents’ growing presence in the software industry([50, 3, 11, 7, 8]). The advancement in AI agents is rapid, highlighted by the AI assistant tools like MultiOn ([9]) and Adept AI([25]), and AI consumer products like Rabbit R1 ([26]) and Humane AI Pin([5]), which are gaining traction in the consumer sector. Research into AI agents has been robust, witnessing developments in the chain of thought ([58, 16]) and enhanced prompting techniques([51]). Moreover, the rise of multi-agent systems ([53, 43, 39, 29]) marks a novel trend in the industry, showcasing the use of language models to develop dependable software that empowers users([54, 38]). These innovations leverage the API calling ([56, 12, 49]) and reasoning abilities ([40, 36]) of large, cloud-based language models to convert human natural language instructions into actionable commands. Despite the considerable progress in creating valuable AI agents, reliance on cloud models raises issues concerning privacy, inference costs, and the need for Wi-Fi connectivity ([57, 22]).
@@
The cost of using large language models like Google’s Gemini family models ([45]) and OPENAI’s GPT series models ([33, 34, 2, 1]) can be substantial; for example, an hour-long interaction with an AI bot might cost around 0.24 USD according to the GPT-4 API price. When it comes to function calling, employing RAG-based ([17, 27, 19, 15]) or context-augmented ([35]) methods requires processing about 1000 tokens for each call, resulting in costs of approximately 0.01 USD. In practical applications, where hundreds of function calls may be made, the cumulative cost can be much. Additionally, the potential for privacy violations deters many from using GPT-4, amid concerns that sensitive information might be exposed.
@@
To mitigate costs and enhance privacy, there is a trend towards creating smaller models for deployment on edge devices like smartphones, cars, VR headsets, and personal computers ([52, 6, 21, 18, 55, 13, 41]). However, edge computing-based models face challenges with slower latency, far from production readiness, and the limited battery life of edge devices further complicates continuous interaction. Research shows that energy consumption reaches 0.1J per token for 1 billion parameter models ([23]). Therefore, employing a 7B parameter model for function calls with traditional retrieval-augmented methods would consume 700J per call, roughly 1.4% of a 50kJ iPhone battery, limiting to around 71 function calls.
@@
Smaller models often fall short in reasoning tasks and demand extensive tuning for effective function calling. To address these issues, we developed a method to enhance both accuracy and latency for function calling on 2B parameter models on devices, achieving state-of-the-art (SOTA) results. This approach involves tokenizing the core function’s name and fine-tuning the model with functional tokens. Fine-tuning with these tokens allows the model to understand software application capabilities with additional special tokens, learning to map function descriptions to specific tokens. In the inference phase, the model uses functional tokens to achieve better performance in function calling compared to GPT-4. We present a 2B parameter model fine-tuned from Gemma 2B ([10]), saving over 95% context length during model inference. For iPhone use, this enables 37 times more function calls with the same battery and reduces latency by approximately 35 times per function call.
@@ 2 Related works
Deployment of on-device language models  Due to memory limitations and lower inference speeds, deploying larger models on edge devices like PCs or smartphones is challenging. Nonetheless, efforts to deploy smaller-scale Large Language Models (LLMs) to edge devices are underway. Open-source models of manageable sizes, such as Gemma-2B, Gemma-7B, StableCode-3B ([31]), and Llama-7B ([47]), have been introduced. To enhance these models’ inference speed on devices, research initiatives like Llama cpp ([24]) have been developed. The MLC LLM framework ([46]) allows the operation of 7B language models on mobile phones and other edge devices, demonstrating compatibility across various hardware, including AMD, NVIDIA, Apple, and Intel GPUs.
Function calling in language models  Rapid advancements have been observed in the function-calling capabilities of smaller-scale models. Projects such as NexusRaven ([42]), Toolformer ([37]), ToolAlpaca ([44]), Gorrilla ([30]), ToolLlama ([32]) and Taskmatrix ([20]) have demonstrated that 7B and 13B models can call external APIs with efficacy comparable to GPT-4. The pioneering Octopus v1 project even enabled a 2B model to perform on par with GPT-4. This body of work utilizes a RAG-based method for function calling, where the model retrieves relevant functions from a large pool based on the user’s query, then generates a response using these functions as context.
Fine-tuning and adaptors of language models  Fine-tuning language models has become a prevalent practice, with various efforts dedicated to this endeavor. LoRA ([14]) is often the method of choice for training models under GPU resource constraints. We use both full model training and LoRA training in our work, and compare their performance. A notable benefit of LoRA is its facilitation of extended functionalities in models, suggesting its potential to adapt our current framework for a broad range of applications.
@@ 3.1 Causal language model as a classification model
where $q$ denotes the query, $f$ signifies the selected function name, and params represent the parameters for the chosen function. Driven by the principles of multitask learning/meta learning ([4]), and to achieve quicker inference speeds and system convenience, we pursued a unified GPT model strategy, setting $\pi_{1}=\pi_{2}=\pi$. Therefore, we redefine our objective as:
@@ 3.2 Functional token
The critical innovation presented in this paper is the introduction of the functional token. Drawing parallels to the tokenization of natural language, we propose representing specific functions as functional tokens. We introduce a training methodology for these tokens, inspired by techniques used in natural language models for handling rare words. This training approach is inspired by the word2vec ([28]) framework, where contextual words enrich a token’s semantic representation. For instance, pretrained language models may initially struggle to recognize specialized terms such as PEGylation and Endosomal Escape from the domain of chemistry. However, these models can learn such terms through causal language modeling, leveraging corpora that include these specialized terms. Similarly, functional tokens can be learned using comparable methods. Our experiments have not identified any limitations on the number of functional tokens that can be defined, allowing users to map any specific function to a token.
@@ 4.1 Android function calls
Llama-7B RAG Evaluation  Initially, the pretrained Llama-7B model showed limited ability in generating the expected outcomes, leading us to employ a Llama-7B variant fine-tuned for function call generation ([48]). For the Llama-7B assessment, we applied the RAG method without strict output format requirements, considering responses with missing parentheses as correct. This evaluation was conducted on a single NVIDIA A100 machine, with all results from Llama-7B compared against the ground truth. The primary errors were incorrect function name selection and erroneous parameter generation. Despite employing few-shot learning to guide the model towards accurate function generation, the performance was modest, with an accuracy rate of 68.095% when overlooking format requirements and a latency of 13.46 seconds, excluding model loading time. To improve latency, we implemented optimizations such as flash attention and a fast tokenizer.

### v5

@@ 1 Introduction
Large language models have demonstrated impressive capabilities in function calling, significantly contributing to AI agents’ growing presence in the software industry [50, 3, 11, 7, 8]. The advancement in AI agents is rapid, highlighted by the AI assistant tools like MultiOn [9] and Adept AI [25], and AI consumer products like Rabbit R1 [26] and Humane AI Pin [5], which are gaining traction in the consumer sector. Research into AI agents has been robust, witnessing developments in the chain of thought [58, 16] and enhanced prompting techniques [51]. Moreover, the rise of multi-agent systems [53, 43, 39, 29] marks a novel trend in the industry, showcasing the use of language models to develop dependable software that empowers users [54, 38]. These innovations leverage the API calling [56, 12, 49] and reasoning abilities [40, 36] of large, cloud-based language models to convert human natural language instructions into actionable commands. Despite the considerable progress in creating valuable AI agents, reliance on cloud models raises issues concerning privacy, inference costs, and the need for Wi-Fi connectivity [57, 22].
@@
The cost of using large language models like Google’s Gemini family models [45] and OPENAI’s GPT series models [33, 34, 2, 1] can be substantial; for example, an hour-long interaction with an AI bot might cost around 0.24 USD according to the GPT-4 API price. When it comes to function calling, employing RAG-based [17, 27, 19, 15] or context-augmented [35] methods requires processing about 1000 tokens for each call, resulting in costs of approximately 0.01 USD. In practical applications, where hundreds of function calls may be made, the cumulative cost can be much. Additionally, the potential for privacy violations deters many from using GPT-4, amid concerns that sensitive information might be exposed.
@@
To mitigate costs and enhance privacy, there is a trend towards creating smaller models for deployment on edge devices like smartphones, cars, VR headsets, and personal computers [52, 6, 21, 18, 55, 13, 41]. However, edge computing-based models face challenges with slower latency, far from production readiness, and the limited battery life of edge devices further complicates continuous interaction. Research shows that energy consumption reaches 0.1J per token for 1 billion parameter models [23]. Therefore, employing a 7B parameter model for function calls with traditional retrieval-augmented methods would consume 700J per call, roughly 1.4% of a 50kJ iPhone battery, limiting to around 71 function calls.
@@
Smaller models often fall short in reasoning tasks and demand extensive tuning for effective function calling. To address these issues, we developed a method to enhance both accuracy and latency for function calling on 2B parameter models on devices, achieving state-of-the-art (SOTA) results. This approach involves tokenizing the core function’s name and fine-tuning the model with functional tokens. Fine-tuning with these tokens allows the model to understand software application capabilities with additional special tokens, learning to map function descriptions to specific tokens. In the inference phase, the model uses functional tokens to achieve better performance in function calling compared to GPT-4. We present a 2B parameter model fine-tuned from Gemma 2B [10], saving over 95% context length during model inference. For iPhone use, this enables 37 times more function calls with the same battery and reduces latency by approximately 35 times per function call.
@@ 2 Related works
Deployment of on-device language models  Due to memory limitations and lower inference speeds, deploying larger models on edge devices like PCs or smartphones is challenging. Nonetheless, efforts to deploy smaller-scale Large Language Models (LLMs) to edge devices are underway. Open-source models of manageable sizes, such as Gemma-2B, Gemma-7B, StableCode-3B [31], and Llama-7B [47], have been introduced. To enhance these models’ inference speed on devices, research initiatives like Llama cpp [24] have been developed. The MLC LLM framework [46] allows the operation of 7B language models on mobile phones and other edge devices, demonstrating compatibility across various hardware, including AMD, NVIDIA, Apple, and Intel GPUs.
Function calling in language models  Rapid advancements have been observed in the function-calling capabilities of smaller-scale models. Projects such as NexusRaven [42], Toolformer [37], ToolAlpaca [44], Gorrilla [30], ToolLlama [32] and Taskmatrix [20] have demonstrated that 7B and 13B models can call external APIs with efficacy comparable to GPT-4. The pioneering Octopus v1 project even enabled a 2B model to perform on par with GPT-4. This body of work utilizes a RAG-based method for function calling, where the model retrieves relevant functions from a large pool based on the user’s query, then generates a response using these functions as context.
Fine-tuning and adaptors of language models  Fine-tuning language models has become a prevalent practice, with various efforts dedicated to this endeavor. LoRA [14] is often the method of choice for training models under GPU resource constraints. We use both full model training and LoRA training in our work, and compare their performance. A notable benefit of LoRA is its facilitation of extended functionalities in models, suggesting its potential to adapt our current framework for a broad range of applications.
@@ 3.1 Causal language model as a classification model
where $q$ denotes the query, $f$ signifies the selected function name, and params represent the parameters for the chosen function. Driven by the principles of multitask learning/meta learning [4], and to achieve quicker inference speeds and system convenience, we pursued a unified GPT model strategy, setting $\pi_{1}=\pi_{2}=\pi$. Therefore, we redefine our objective as:
@@ 3.2 Functional token
The critical innovation presented in this paper is the introduction of the functional token. Drawing parallels to the tokenization of natural language, we propose representing specific functions as functional tokens. We introduce a training methodology for these tokens, inspired by techniques used in natural language models for handling rare words. This training approach is inspired by the word2vec [28] framework, where contextual words enrich a token’s semantic representation. For instance, pretrained language models may initially struggle to recognize specialized terms such as PEGylation and Endosomal Escape from the domain of chemistry. However, these models can learn such terms through causal language modeling, leveraging corpora that include these specialized terms. Similarly, functional tokens can be learned using comparable methods. Our experiments have not identified any limitations on the number of functional tokens that can be defined, allowing users to map any specific function to a token.
@@ 4.1 Android function calls
Llama-7B RAG Evaluation  Initially, the pretrained Llama-7B model showed limited ability in generating the expected outcomes, leading us to employ a Llama-7B variant fine-tuned for function call generation [48]. For the Llama-7B assessment, we applied the RAG method without strict output format requirements, considering responses with missing parentheses as correct. This evaluation was conducted on a single NVIDIA A100 machine, with all results from Llama-7B compared against the ground truth. The primary errors were incorrect function name selection and erroneous parameter generation. Despite employing few-shot learning to guide the model towards accurate function generation, the performance was modest, with an accuracy rate of 68.095% when overlooking format requirements and a latency of 13.46 seconds, excluding model loading time. To improve latency, we implemented optimizations such as flash attention and a fast tokenizer.